OpenStudy (anonymous):

factorise cos x + cos 3x + cos 5x + cos 7x

OpenStudy (gorv):

CosC + CosD = 2Cos[(C+D)/2]Cos[(C-D)/2] use this formula

OpenStudy (gorv):

cos x + cos 3x=2Cos[(x+3x)/2]Cos[(x-3x)/2]=2*cos2x*cosx cos 5x + cos 7x=2Cos[(5x+7x)/2]Cos[(5x-7x)/2]=2*cos6x*cosx add these cos x + cos 3x + cos 5x + cos 7x=2*cos2x*cosx+2*cos6x*cosx =2*cosx(cos2x+cos6x) again apply C & D formula 2*cosx(cos2x+cos6x)=2*cosx*(2Cos[(2x+6x)/2]Cos[(2x-6x)/2]) =2*cosx(cos4x*cos2x) =2*cosx*cos2x*cos4x